OpenStudy (anonymous):

The speed of the current in a river is 6 mph. A ferry operator who works that part of the river is looking to buy a new boat for his business. Every day, his route takes him 22.5 miles each way against the current and back to his dock, and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator’s needs?

OpenStudy (anonymous):

22.5/(x+6) + 22.5/(x-6) = 9

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

I still don't understand ......

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

hold on while i get someone

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

i don't know sorry. that is as much as i could help! Ask an ambassador for help.

jigglypuff314 (jigglypuff314):

if river current pushes 6mph ---> that way then when you're going --> you'll be pushed by the river while going <-- you'll be pulled if you let your speed be represented by " x " then your speed going --> way would be x+6 and your speed going <-- way would be x-6 follow so far? :)

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

then distance divided by speed = time so since distance = 22.5miles twice (back and forth) and you have different speeds going back and forth 22.5 22.5 ----- + ----- = 9 (x+6) (x-6)

OpenStudy (anonymous):

ok so where do we go from there?

jigglypuff314 (jigglypuff314):

solve for x... I multiplied both sides by (x+6)(x-6) to get 22.5(x-6) + 22.5(x+6) = 9(x+6)(x-6) then use distributive property

OpenStudy (anonymous):

9(b^2-36) = 9b^2 - 324

jigglypuff314 (jigglypuff314):

mmm yep 22.5(x-6) + 22.5(x+6) = 9(x+6)(x-6) 22.5x - 135 + 22.5x + 135 = 9x^2 - 324 -> 9x^2 - 45x - 324 = 0 (solve for x)

OpenStudy (anonymous):

where did the minus 45x come from?

jigglypuff314 (jigglypuff314):

22.5x + 22.5x = 45x

jigglypuff314 (jigglypuff314):

mmm if you were confused...? " 22.5(x-6) + 22.5(x+6) = 9(x+6)(x-6) " means 9(x+6)(x-6) - 22.5(x-6) - 22.5(x+6) which I simplified to 9x^2 - 45x - 324

jigglypuff314 (jigglypuff314):

then factor out a 9 then factor... what do you get?